# Global bifurcation of localised 2D patterns emerging from spatial heterogeneity

### v1

@@ 3.2 Global bifurcation
We now prove that the local bifurcation curves found in the previous section can be continued away from the bifurcation point; we call these global bifurcation branches, and use the analytic global implicit function theorem of Chen, Walsh and Wheeler [17, Theorem B.1] (which we present as Theorem A.3 in the appendix) in order to prove their existence.
@@ 4. IV:
This result follows directly from Theorem A.3 with $X=H^{4}_{\mathrm{e}}$ and $Y=L^{2}_{\mathrm{e}}$. While the bifurcation point $(\varepsilon_{k,n},0)$ does not satisfy the conditions for Theorem A.3, since $\mathrm{d}_{2}\mathscr{F}[\varepsilon_{k,n},0]$ is not an isomorphism $H^{4}_{\mathrm{e}}\to L^{2}_{\mathrm{e}}$, it follows that there is a point $(\varepsilon_{*},v_{*})\in\mathcal{C}_{\delta}$ for which Theorem A.3 holds. This is due to the Fredholm properties of $\mathcal{L}_{\phi}:=\mathrm{d}_{2}\mathscr{F}[\varepsilon,\phi]$ proved in Lemma 2.1 for any $(\varepsilon,\phi)\in U$, as well as the transversality property proved in Lemma 3.2. The possibility of a loss of Fredholm properties in Theorem A.3 is also avoided since we proved that $\mathrm{d}_{2}\mathscr{F}[\varepsilon,\phi]$ is Fredholm for all $\varepsilon>0$, $\phi\in H^{4}_{\mathrm{e}}$ in Lemma 2.1. ∎
@@
    We note that $\mathcal{C}$ having an analytic parametrisation at each point does not imply that $\mathcal{C}$ is smooth. In fact, as discussed in [11, Remark 9.1.2 (3)], the curve $\mathcal{C}$ can possess isolated points where $\mathcal{C}$ is not even continuously differentiable. We demonstrate an example where $\mathcal{C}$ has an isolated corner in Figure 8(iv).
@@
    Different versions of the alternatives I-III are presented in the literature (see, for example [13, Remarks 4.2 and 4.3]); in particular, Constantin, Strauss and Vărvărucă [18, Remark 8] highlight that the alternatives presented by Buffoni and Toland [11, Theorem 9.1.1] are not exhaustive, and present a different formulation in their global bifurcation theorem [18, Theorem 6]. Here we use the same formulation of the alternatives as used by Chen, Walsh and Wheeler [17, Theorem B.1], which we present in more detail in Theorem A.3. Note that alternatives $I$ and $II$ are often indistinguishable from one another and so we combine the two into alternative $(i)$ in Theorem A.3.
@@
    In many applications one can exclude the case when $\mathcal{C}$ forms a closed loop in part $(c)$, typically via bifurcation theory on cones [11, Theorem 9.2.2] and using nodal properties of the bifurcating solution. However, in our problem we numerically observe cases where $\mathcal{C}$ forms a closed loop (such as in Figure 8(iv)) and so we do not expect this to be possible in general.
@@
    The loss of compactness alternative can sometimes be excluded through certain properties of solutions; for example, Haziot and Wheeler [32] used the monotonicity of their solutions while Kozlov [49] utilised uniform decay rates. We have been unable to exclude a loss of compactness in the current setting, but it may be possible to further strengthen our global bifurcation result in future studies.
@@ 4 Discussion
This work lays the foundation for further exploration of localised patterns induced by compact spatial heterogeneities. A straightforward extension would be to consider the three-dimensional analogue of (1.2), where the Fredholm properties will remain the same and the point spectrum can again be computed explicitly using spherical harmonics and spherical Bessel functions. Replacing the step function $V$ by a smooth potential would be a more realistic model for spatial heterogeneities in experiments, and one would expect much of the underlying theory of this work to persist in that case. In particular, we highlight the approach presented by Brooks, Derks and Lloyd [9, Section 5] in which they prove the persistence of fronts in the sine-Gordon equation from piecewise to smooth heterogeneities; we expect a similar argument to hold in this work using the approach of van Heijster and Sandstede [62].
@@ Acknowledgments
The authors are grateful to Björn de Rijk, Mark D. Groves and Miles Wheeler for their assistance on certain aspects of this manuscript. DJH acknowledges support from the Alexander von Humboldt Foundation, and MT and DJBL acknowledge support from EPRSC grant UKRI070. The authors also thank Tom Bridges, Ryan Goh, Bastian Hilder and Jörg Weber for helpful discussions and suggestions regarding this work. For the purpose of Open Access, the authors have applied a Creative Commons Attribution (CC BY) public copyright licence to any Author Accepted Manuscript version arising from this submission.
@@ Appendix A Bifurcation theorems
In our local bifurcation theorem (Theorem 3.3) we use an analytic version of the Crandall–Rabinowitz theorem (cf. Buffoni and Toland [11, Theorem 8.3.1]), which we state here with the notation presented by Groves and Horn [30, Theorem 3.1].
@@
A proof of the standard Crandall–Rabinowitz theorem can be found in the original paper by Crandall and Rabinowitz [19, Theorem 1], while the analytic version is proved by Buffoni and Toland [11, Theorem 8.3.1]. Using the analytic properties of the local bifurcation curve, we can also apply the following supplementary theorem presented by Groves and Horn [30, Theorem 4.1] to characterise the local bifurcation curve $\mathcal{C}_{\delta}$.
@@
This theorem uses the analytic properties of bifurcating solutions in order to take Taylor series expansions for $|s|\ll 1$ and characterise the bifurcation structure by standard normal forms in bifurcation theory (see, for example Verhulst [63, Section 13.5]).
To prove our global bifurcation result (Theorem 3.9) we use the analytic global implicit function theorem given by Chen, Walsh and Wheeler [17, Theorem B.1], which extends the global bifurcation theorem of Buffoni and Toland [11, Theorem 9.1.1] to cases where compactness and Fredholm properties of solutions are no longer assumed.
@@
For a proof of this theorem, we refer the reader to Chen, Walsh and Wheeler [17, Theorem B.1], along with Buffoni and Toland [11, Theorem 9.1.1] for additional details.
@@ Appendix B Additional proofs
We begin with a proof that multiplication by an element in $H^{m}=H^{m}(\mathbb{R}^{2})$ is a compact operator from $H^{m}$ to $L^{2}=L^{2}(\mathbb{R}^{2})$ with $m>1$; while this result is often used in linear stability analysis (cf. Kapitula and Promislow [46, Theorem 3.1.11]), we present a proof here for completeness.

### v2

@@ 3.2 Global bifurcation
We now prove that the local bifurcation curves found in the previous section can be continued away from the bifurcation point; we call these global bifurcation branches, and use the analytic global implicit function theorem of Chen, Walsh and Wheeler [19, Theorem B.1] (which we present as Theorem A.3 in the appendix) in order to prove their existence.
@@ 4. IV:
This result follows directly from Theorem A.3 with $X=H^{4}_{\mathrm{e}}$ and $Y=L^{2}_{\mathrm{e}}$. Upon choosing some $(\varepsilon_{*},u_{*})\in\mathcal{C}_{\delta}$ such that the linearisation of (1.3) at $(\varepsilon_{*},u_{*})$ is an isomorphism, one can apply the analytic implicit function theorem and use theory regarding analytic curves to globally extend the local solution branch; see, for example, the proofs of Theorem 9.1.1 in [14], Theorem 6.1 in [77], or Theorem 2.4 in [24]. One typically continues the local branch $\mathcal{C}_{\delta}$ from a starting point $(\varepsilon_{*},u_{*})\in\mathcal{C}_{\delta}$ which is distinct from the point $(\varepsilon_{0},0)$; this is because either $(\varepsilon_{0},0)$ is a bifurcation point and thus the linear operator is not invertible at $\varepsilon=\varepsilon_{0}$, such as in [14, Proof of Theorem 9.1.1], or $(\varepsilon_{0},0)$ lies on the boundary of the set $U$, such as in [77, Proof of Theorem 6.1].
By Lemma 2.1, $\mathcal{L}_{\phi}:=\mathrm{d}_{2}\mathscr{F}[\varepsilon,\phi]$ is a Fredholm operator with index zero along the local bifurcation branch $\mathcal{C}_{\delta}$ and by Lemma 3.2 an isolated eigenvalue of $\mathcal{L}_{\phi}$ passes through zero at $(\varepsilon,\phi)=(\varepsilon_{0},0)$. It follows that there is a non-empty subset $\mathcal{C}_{*}\subseteq\mathcal{C}_{\delta}$ along which $\mathcal{L}_{\phi}$ is an isomorphism; we choose some point $(\varepsilon_{*},u_{*})\in\mathcal{C}_{*}$ and apply Theorem A.3.
@@
    We note that $\mathcal{C}$ having an analytic parametrisation at each point does not imply that $\mathcal{C}$ is smooth. In fact, as discussed in [14, Remark 9.1.2 (3)], the curve $\mathcal{C}$ can possess isolated points where $\mathcal{C}$ is not even continuously differentiable. We demonstrate an example where $\mathcal{C}$ has an isolated corner in Figure 8(iv).
@@
    Different versions of the alternatives I-III are presented in the literature (see, for example [9, Remarks 4.2 and 4.3]); in particular, Constantin, Strauss and Vărvărucă [20, Remark 8] highlight that the alternatives presented by Buffoni and Toland [14, Theorem 9.1.1] are not exhaustive, and present a different formulation in their global bifurcation theorem [20, Theorem 6]. Here we use the same formulation of the alternatives as used by Chen, Walsh and Wheeler [19, Theorem B.1], which we present in more detail in Theorem A.3. Note that alternatives $I$ and $II$ are often indistinguishable from one another and so we combine the two into alternative $(i)$ in Theorem A.3.
@@
    In many applications one can exclude the case when $\mathcal{C}$ forms a closed loop in part $(c)$, typically via bifurcation theory on cones [14, Theorem 9.2.2] and using nodal properties of the bifurcating solution. However, in our problem we numerically observe cases where $\mathcal{C}$ forms a closed loop (such as in Figure 8(iv)) and so we do not expect this to be possible in general.
@@
    The loss of compactness alternative can sometimes be excluded through certain properties of solutions; for example, Haziot and Wheeler [35] used the monotonicity of their solutions while Kozlov [53] utilised uniform decay rates. For particular choices of nonlinear function $f$ where one can prove that there are no non-trivial solutions to the homogeneous problem (i.e. with $V\equiv 0$), such as $f(u)=-u^{3}$, one might be able to apply a concentration-compactness argument [54, 55] to eliminate the loss of compactness. We have been unable to exclude a loss of compactness in the current setting outside of axisymmetric patterns—which we present in the following lemma and discuss further in Section B.1 of the Appendix—but it may be possible to further strengthen our global bifurcation result in future studies.
In the case of axisymmetric solutions $u\in X^{4}_{0}$, we obtain the following pre-compactness result.
Lemma 3.11.
Let $\mathcal{C}\subset X^{4}_{0}$ be a global bifurcation curve for a localised axisymmetric solution to (1.3). Then, alternative IV in Theorem 3.9 cannot occur.
We prove a more general version of this result (Lemma B.2) in the Appendix, where we rely on an a priori uniform decay rate for radial functions. Such a uniform decay rate is not known for non-radial functions, and we would not expect there to be an analogous result for all functions due to translational invariance of the space $H^{4}(\mathbb{R}^{2})$; for further discussion of alternative $IV$, see Section B.1.
@@
We briefly comment on the boundary $\varepsilon=0$; this corresponds to the parameter values at which the essential spectrum of the linear operator $\mathcal{L}$ touches the imaginary axis, known as the onset of essential instability [69]. Essential instabilities are typical of homogeneous pattern-forming systems, such as the Swift-Hohenberg equation (1.2) with $V\equiv 0$, where Crandall-Rabinowitz bifurcation theory can no longer be applied due to the loss of Fredholm properties at $\varepsilon=0$. There have been several studies of localised states at the onset of essential instability in one spatial dimension, using techniques from spatial dynamics [71, 70], but extending such analysis to higher spatial dimensions remains an open problem.
@@ 4 Discussion
This work lays the foundation for further exploration of localised patterns induced by compact spatial heterogeneities. A straightforward extension would be to consider the three-dimensional analogue of (1.2), where the Fredholm properties will remain the same and the point spectrum can again be computed explicitly using spherical harmonics and spherical Bessel functions. Replacing the step function $V$ by a smooth potential would be a more realistic model for spatial heterogeneities in experiments, and one would expect much of the underlying theory of this work to persist in that case. In particular, we highlight the approach presented by Brooks, Derks and Lloyd [12, Section 5] in which they prove the persistence of fronts in the sine-Gordon equation from piecewise to smooth heterogeneities; we expect a similar argument to hold in this work using the approach of van Heijster and Sandstede [75].
@@ Acknowledgments
The authors are grateful to Björn de Rijk, Mark D. Groves and Miles Wheeler for their assistance on certain aspects of this manuscript. DJH acknowledges support from the Alexander von Humboldt Foundation, and MT and DJBL acknowledge support from EPRSC grant UKRI070. The authors also thank Tom Bridges, Ryan Goh, Bastian Hilder, Jörg Weber, Gui-Qiang Chen, Melanie Rupflin and Sergey Zelik for helpful discussions and suggestions regarding this work. For the purpose of Open Access, the authors have applied a Creative Commons Attribution (CC BY) public copyright licence to any Author Accepted Manuscript version arising from this submission.
@@ Appendix A Bifurcation theorems
In our local bifurcation theorem (Theorem 3.3) we use an analytic version of the Crandall–Rabinowitz theorem (cf. Buffoni and Toland [14, Theorem 8.3.1]), which we state here with the notation presented by Groves and Horn [32, Theorem 3.1].
@@
A proof of the standard Crandall–Rabinowitz theorem can be found in the original paper by Crandall and Rabinowitz [21, Theorem 1], while the analytic version is proved by Buffoni and Toland [14, Theorem 8.3.1]. Using the analytic properties of the local bifurcation curve, we can also apply the following supplementary theorem presented by Groves and Horn [32, Theorem 4.1] to characterise the local bifurcation curve $\mathcal{C}_{\delta}$.
@@
This theorem uses the analytic properties of bifurcating solutions in order to take Taylor series expansions for $|s|\ll 1$ and characterise the bifurcation structure by standard normal forms in bifurcation theory (see, for example Verhulst [76, Section 13.5]).
To prove our global bifurcation result (Theorem 3.9) we use the analytic global implicit function theorem given by Chen, Walsh and Wheeler [19, Theorem B.1], which extends the global bifurcation theorem of Buffoni and Toland [14, Theorem 9.1.1] to cases where compactness and Fredholm properties of solutions are no longer assumed.
@@
For a proof of this theorem, we refer the reader to Chen, Walsh and Wheeler [19, Theorem B.1], along with Buffoni and Toland [14, Theorem 9.1.1] for additional details.
@@ Appendix B Additional proofs
We begin with a proof that multiplication by an element in $H^{m}=H^{m}(\mathbb{R}^{2})$ is a compact operator from $H^{m}$ to $L^{2}=L^{2}(\mathbb{R}^{2})$ with $m>1$; while this result is often used in linear stability analysis (cf. Kapitula and Promislow [50, Theorem 3.1.11]), we present a proof here for completeness.
@@
B.1 Pre-compactness of solution branches
In this section we attempt to prove that alternative $IV$ in Theorem 3.9 does not occur; that is, branches of solutions to (1.3) do not lose compactness for $\varepsilon>0$. For the remainder of this section we consider the following equation
$\mathcal{L}_{\varepsilon}\,u=V_{\varepsilon}\,u+f_{\varepsilon}(u),$ (B.1)
where $\varepsilon>0$ is a fixed parameter, $u\in H^{2m}(\mathbb{R}^{d})$ with $m>d/4$, and $f_{\varepsilon}$ is some real-analytic function that satisfies $f_{\varepsilon}(0)=f_{\varepsilon}^{\prime}(0)=0$ and may depend on the parameter $\varepsilon$. The linear operator $\mathcal{L}_{\varepsilon}:H^{2m}(\mathbb{R}^{d})\to L^{2}(\mathbb{R}^{d})$ and potential function $V_{\varepsilon}=V_{\varepsilon}(\mathbf{x})$ are assumed to satisfy the following properties:
  • $\mathcal{L}_{\varepsilon}:H^{2m}(\mathbb{R}^{d})\to L^{2}(\mathbb{R}^{d})$ is defined as a Fourier multiplier, such that $\mathcal{F}[\mathcal{L}_{\varepsilon}u](\mathbf{k})=\ell_{\varepsilon}(\mathbf{k})\,\mathcal{F}[u](\mathbf{k})$, where there exists some $d_{\varepsilon}>0$ such that the Fourier symbols $\ell_{\varepsilon}$ satisfies
    $\ell_{\varepsilon}(\mathbf{k})\geq d_{\varepsilon}\quad\text{for all $\mathbf{k}\in\mathbb{R}^{d}$,}\qquad\qquad|\ell_{\varepsilon}(\mathbf{k})|=\mathcal{O}(|\mathbf{k}|^{2m})\quad\text{as $|\mathbf{k}|\to\infty$.}$
  • $V_{\varepsilon}\in L^{\infty}(\mathbb{R}^{d})$ and satisfies the following decay condition; for any $\delta>0$ there is some $T>0$ such that
    $|V_{\varepsilon}(\mathbf{x})|\leq\delta\qquad\text{for all $|\mathbf{x}|>T$.}$
We emphasise that both the Swift-Hohenberg equation (1.2) and the more general PDE (1.1) satisfy the above assumptions for $\mathcal{L}_{\varepsilon}$ and $V_{\varepsilon}$, and so we can apply any results for (B.1) directly to (1.3).
We now prove that the global solution curve of an axisymmetric solution to (B.1) does not lose compactness. Note that in the following proof, with slight abuse of notation, we use $X^{2m}_{0}$ to denote the subspace of radially symmetric functions in $H^{2m}(\mathbb{R}^{d})$.
Lemma B.2.
Let $\mathcal{C}\subset X^{2m}_{0}$ be a global bifurcation curve for a localised axisymmetric solution to (B.1). Then, alternative IV in Theorem 3.9 cannot occur.
Proof.
Let $\{u_{n}\}$ be a bounded sequence in $X^{2m}_{0}$ of solutions to (B.1) and hence weakly convergent to some $u\in X^{2m}_{0}$. We note that $u_{n}\to u$ (up to subsequences) in $L^{2}(B_{T})$ for any $T>0$, where $B_{T}$ denotes the unit ball of radius $T$ about the origin.
By Strauss’ radial lemma [72, Lemma 1], there exists a constant $c_{1}>0$ such that
$|u(\mathbf{x})|\leq c_{1}\,|\mathbf{x}|^{-\frac{(d-1)}{2}}\,\|u\|_{H^{1}(\mathbb{R}^{d})}$ (B.2)
holds for all $u\in X^{1}_{0}$. The condition that $f_{\varepsilon}$ satisfies $f_{\varepsilon}(0)=f_{\varepsilon}^{\prime}(0)=0$ also implies that
$|f_{\varepsilon}(u)|\leq c_{2}|u|^{\alpha+1}$
for all $u\in H^{2m}(\mathbb{R}^{d})$ and some $\alpha>0$, where the constant $c_{2}>0$ is uniformly bounded in $H^{2m}(\mathbb{R}^{d})$ (since $\|u\|_{L^{\infty}}$ is uniformly bounded). In particular, coupled with the uniform decay estimate (B.2), it follows that
$|f_{\varepsilon}(u(\mathbf{x}))|\leq c_{2}|u(\mathbf{x})|^{\alpha}\,|u(\mathbf{x})|\leq c_{1}^{\alpha}\,c_{2}\,\|u\|_{H^{1}(\mathbb{R}^{d})}^{\alpha}T^{-\frac{(d-1)\alpha}{2}}|u(\mathbf{x})|,\qquad\qquad\text{for all $|\mathbf{x}|\geq T$}$
and so
$\|f_{\varepsilon}(u)\|^{2}_{L^{2}(\mathbb{R}^{d}\backslash B_{T})}\leq c_{1}^{2\alpha}\,c_{2}^{2}\,\|u\|_{H^{1}(\mathbb{R}^{d})}^{2\alpha}\|u\|^{2}_{L^{2}(\mathbb{R}^{d})}\,T^{-(d-1)\alpha}.$
To prove that $u_{n}\to u$ strongly in $H^{2m}(\mathbb{R}^{d})$, we note that $\mathcal{L}_{\varepsilon}:H^{2m}(\mathbb{R}^{d})\to L^{2}(\mathbb{R}^{d})$ is an isomorphism with
$\|u\|_{H^{2m}(\mathbb{R}^{d})}\leq\sup_{\mathbf{k}\in\mathbb{R}^{d}}\left(\frac{(1+|\mathbf{k}|^{2})^{m}}{\ell_{\varepsilon}(\mathbf{k})}\right)\,\|\mathcal{L}_{\varepsilon}u\|_{L^{2}(\mathbb{R}^{d})}=:C_{\varepsilon}\,\|\mathcal{L}_{\varepsilon}u\|_{L^{2}(\mathbb{R}^{d})}$
for all $u\in H^{2m}(\mathbb{R}^{d})$. It thus follows that
[…]
for any $\kappa>0$. The final step follows by choosing $T$ sufficiently large so that $T^{-(d-1)\alpha}$ and $\delta^{2}$ are both arbitrarily small, and then choosing $N=N(T)>0$ so that the first two terms are arbitrarily small for all $n>N$ (where we use that the each coefficient is uniformly bounded in $H^{2m}(\mathbb{R}^{d})$, $u_{n}\to u$ strongly in $L^{2}(B_{T})$ for any $T>0$, and $f_{\varepsilon}$ is at least continuous). ∎
The main difficulty in extending Lemma B.2 beyond axisymmetric solutions lies in reproducing the uniform decay estimate provided by Strauss’ radial lemma. Categorising the uniform decay in general for $u\in H^{1}(\mathbb{R}^{d})$ is significantly more difficult, as such functions are invariant under spatial translations and so we cannot expect that an estimate of the form (B.2) holds for all $u\in H^{1}(\mathbb{R}^{d})$.
We present a slightly different formulation of Strauss’ radial lemma for the special case $d=2$, which now holds for all $u\in H^{1}(\mathbb{R}^{2})$.
Lemma B.3.
Fix $k\in\mathbb{Z}$, suppose $u\in H^{1}(\mathbb{R}^{2})$, and define $u_{k}\in H^{1}(\mathbb{R}^{2})$ to be the mode-$k$ component of $u$, given by
$u_{k}(\mathbf{x})=\frac{1}{2\pi}\int_{0}^{2\pi}u(\mathcal{R}_{\theta}\mathbf{x})\,\mathrm{e}^{-\mathrm{i}k\theta}\,\mathrm{d}\theta,$
where $\mathcal{R}_{\theta}$ denotes the standard rotation matrix in $\mathbb{R}^{2}$ about an angle $\theta\in[0,2\pi)$. Then, there exists a constant $C>0$ such that the inequalities
$\begin{split}|u_{k}(\mathbf{x})|\leq{}&(2\pi)^{-\frac{1}{2}}\,C\,|\mathbf{x}|^{-\frac{1}{2}}\,\|u_{k}\|_{H^{1}(\mathbb{R}^{2})},\hskip 18.49988pt\hskip 18.49988pt\int_{0}^{2\pi}|u(\mathcal{R}_{\theta}\mathbf{x})|^{2}\,\mathrm{d}\theta\leq C^{2}\,|\mathbf{x}|^{-1}\,\|u\|^{2}_{H^{1}(\mathbb{R}^{2})},\end{split}$
hold for all $\mathbf{x}\in\mathbb{R}^{2}$. In particular, the constant $C$ is given by $C=\left(\max_{s\geq 0}\{s\,I_{0}(s)\,K_{0}(s)\}\right)^{\frac{1}{2}}$.
Proof.
We begin by noting that any $u\in H^{1}(\mathbb{R}^{2})$ can be expressed via the modal decomposition
$u=\sum_{k\in\mathbb{Z}}u_{k},$ (B.3)
where the infinite sum converges in $H^{1}(\mathbb{R}^{2})$ such that we can write
$\sum_{k\in\mathbb{Z}}\|u_{k}\|_{H^{1}(\mathbb{R}^{2})}^{2}=\|u\|_{H^{1}(\mathbb{R}^{2})}^{2}.$
We note that the mode-$k$ function $u_{k}$ satisfies the following identities
$u_{k}(r\cos\theta,r\sin\theta)=\mathrm{e}^{\mathrm{i}k\theta}\hat{u}_{k}(r),\qquad\qquad\mathcal{F}[u_{k}](\rho\cos\omega,\rho\sin\omega)=\mathrm{i}^{-k}\mathrm{e}^{\mathrm{i}k\omega}\mathcal{H}_{k}[\hat{u}_{k}](\rho),$
where $\hat{u}_{k}$ is termed the radial coefficient of $u_{k}$ (cf. Groves and Hill [33]), and $\mathcal{H}_{k}$ denotes the $k$-index Hankel transform given by
$\mathcal{H}_{k}[\hat{u}_{k}](\rho)=\int_{0}^{\infty}J_{k}(\rho\,r)\,\hat{u}_{k}(r)\,r\,\mathrm{d}r.$
The Hankel transform is self-inverse, and so we obtain the following estimate
[…]
We note that the modal decomposition (B.3) yields the following Fourier expansion for $u(\mathcal{R}_{\theta}\,\cdot)$
$u(\mathcal{R}_{\theta}\,\cdot)=\sum_{k\in\mathbb{Z}}\mathrm{e}^{\mathrm{i}k\theta}\,u_{k},$
and so applying Parseval’s identity yields
$\frac{1}{2\pi}\int_{0}^{2\pi}|u(\mathcal{R}_{\theta}\mathbf{x})|^{2}\,\mathrm{d}\theta=\sum_{k\in\mathbb{Z}}|u_{k}(\mathbf{x})|^{2}.$
Applying the Strauss-type estimate (B.4) for each $k\in\mathbb{Z}$, we obtain
$\int_{0}^{2\pi}|u(\mathcal{R}_{\theta}\mathbf{x})|^{2}\,\mathrm{d}\theta\leq\sum_{k\in\mathbb{Z}}C^{2}\,|\mathbf{x}|^{-1}\|u_{k}\|^{2}_{H^{1}(\mathbb{R}^{2})}=C^{2}\,|\mathbf{x}|^{-1}\|u\|^{2}_{H^{1}(\mathbb{R}^{2})}$
where the last step follows from the $H^{1}$-convergence of (B.3) (which implies that the infinite sum in Parseval’s identity converges uniformly, due to the Weierstrass M-test). ∎
While this lemma provides an equivalent decay rate to that found by Strauss in the axisymmetric case, the result does not provide a uniform decay estimate for $u\in H^{1}(\mathbb{R}^{2})$ but rather on the spherical mean of $u$. This avoids the problem of translational invariance in $H^{1}(\mathbb{R}^{2})$, but means that we are unable to simply apply the same argument as in the proof of Lemma B.2. In particular, we require an estimate of the form
$\int_{0}^{2\pi}|u(\mathcal{R}_{\theta}\mathbf{x})|^{2(1+\alpha)}\,\mathrm{d}\theta\leq c_{1}^{2}\,\rho(T)\,\int_{0}^{2\pi}|u(\mathcal{R}_{\theta}\mathbf{x})|^{2}\,\mathrm{d}\theta\qquad\text{for all $|\mathbf{x}|>T$,}$
where $\rho$ is some function that satisfies $\rho(T)\to 0$ as $T\to\infty$. If such an estimate can be obtained, then the pre-compactness of the global branches would follow identically to the axisymmetric case presented in Lemma B.2.
Remark B.4.
We expect that all bifurcating solutions $u\in H^{2m}(\mathbb{R}^{d})$ to (B.1) decay at an exponential rate as $|\mathbf{x}|\to\infty$. One can characterise this decay property via radial exponential dichotomies (as recently developed by Beck, Goh & Haslam-Hyde [7]) or by studying the associated Green’s function of $\mathcal{L}_{\varepsilon}$ (where one can use Paley–Wiener theory to characterise the decay of a general Fourier symbol $\ell_{\varepsilon}$, such as in the work of Arnesen [1] for one-dimensional Green’s functions). However, in both of these approaches it remains difficult to obtain the uniform exponential decay rates required in order to prove compactness.